Convert log4.5 in the terms of log2 and log3
Answer by lwsshak3(11628)

Convert log4.5 in the terms of log2 and log3
log(4.5)=log(9/2)=log(9)-log(2)=log(3^2)-log(2)=2log(3)-log(2)